How many gallons of 80% antifreeze should be mixed with 15 gallons of 40% antifreeze to obtain a 55% antifreeze mixture?
Answer by mananth(16946)

--------percent ---------------- quantity
Antifreeze I 40 ---------------- 15 gallons
Antifreeze II 80 ---------------- x gallons
Mixture 55 ---------------- 15 + x gallons

40*15+80x=55(15+x)
600+80x=25+55x
80x+-55 x= 825 - -600
25x=225
/ 25
x = 9
9 gallons of Antifreeze II